# Collatz-Weyl Generators: High Quality and High Throughput Parameterized Pseudorandom Number Generators

### v2

@@ 7.2 Results
Table 1 shows that the Collatz-Weyl Generators passed all the single-stream tests, and this confirms their excellent statistical quality. Because the minimum guaranteed period of the CWG32 and CWG64-32 generators was $2^{37}$ and $2^{38}$ bits while they passed the tests up to $2^{43}$ and $2^{44}$ bits, they can generate numbers up to a threshold that is 64 times higher than is required by the minimum guaranteed period without incurring failure. The period of the CWG16 and CWG32-16 generators was $2^{20}$ and $2^{21}$ bits while they passed the tests up to $2^{28}$ and $2^{24}$ bits. Therefore, they can generate numbers up to a threshold 8-256 times higher than is required by the minimum guaranteed period without incurring failure. Hence, we expect that larger-scale versions of these generators would also pass the tests above the threshold of the minimum guaranteed period. NIST, Dieharder, and TestU01 were not applied to the 16-bit versions because they generated too small a number of bits to pass these tests.
@@ 7.3 Birthday Test
Additionally, we performed 1,000 runs of the birthday test on a scaled-down version of the Collatz-Weyl Generator CWG32 and 32768 birthday tests on version CWG16 birthdaytest . First we calculated how many outputs to look at and how many repeats we expect to see, then we collected that many outputs, and counted how many repeats we saw. Finally we calculated p-values for the obtained results. A birthday test can cause some popular PRNGs to quickly fail because non-zero repetitions are expected in a true random number generator over its period of operation. On the contrary, the numbers generated by most PRNGs cannot be repeated in the range of their period. This is not the case with Collatz-Weyl Generators, as explained in the previous sections. The average p-value of CWG32 version in the 1000 runs of the birthday test was $0.507$. In the case of CWG16 we tested all possible $2^{15}$ seeds (every odd $s$) in the full range of minimum guaranteed period ($2^{16}$ outputs). In an ideal random number generator the average expected number of duplicates in range of $2^{16}$ 16-bit numbers is 24109.2, while the average number of duplicates in all streams in CWG16 generator was 24108.4, and the average p-vaule $0.500$. These results are therefore indistinguishable from true random number generation under the birthday test and confirms that the Collatz-Weyl Generators may be used in the full range of minimum guaranteed period without incurring the risk of failure.
@@ 8 COMPARATIVE TESTS OF THROUGHPUT
We measured the time taken by the generators to generate $10^{9}$ random numbers by using an AMD Ryzen 5 4600H, 3 GHz processor with an Ubuntu Windows Subsystem for Linux (WSL) with gcc version 9.4.0. We also assessed the performance of several popular PRNGs for the sake of comparison. Ranlux++ and MixMax17 were run on a MacBook Pro 2.6 GHz, Intel Core i7 CPU in the ROOT library developed by CERN ROOT . Table 3 lists the results.
@@ 9 CONCLUSION AND FUTURE RESEARCH
Let us consider the simplest Collatz-Weyl Generator with $m=2$ that is based on the original Collatz sequences, as shown in Figure 13 (or, equivalently, presented in Figure 14). This generator outputs only one bit per call, but it becomes very difficult for an attacker who does not know $s$, $weyl$, and other bits of $x$ to predict the next or past outputs of the generator after skipping the first few outputs. This is not due only to the non-invertibility of the mapping itself (to iterate back, we have to guess at least one bit at every step even if we know all the states of the generator). The least significant bits of such sequences generated by the proposed algorithm exhibit deterministic chaos, similar to that observed in Collatz sequences and their generalizations kontorovich2010stochastic ; lagarias2023ultimate ; xu2019pseudo ; givens2006conway ; lagarias20103x+ , and thus passed all stringent tests of randomness. This raises the prospect of the possibility of constructing efficient cryptographic primitives based on such schemes—for example, an ultra-lightweight stream cipher. Note that if we define the states of the proposed generator on registers of unlimited size, it seems we obtain a generator with an infinite period. However, we set aside this issue as a hypothesis.
@@
Acknowledgement
I would like to thank professor Dan Tamir, with whom I established collaboration in 2021 and whose article xu2019pseudo inspired me to develop my ideas. During the two years of our collaboration, the schemes I worked on evolved and were significantly modified by me under the pressure of Dan’s critical comments.

### v4

@@ 7.2 Results
Table 1 shows that the Collatz-Weyl Generators passed all the single-stream tests, and this confirms their excellent statistical quality. Because the minimum guaranteed period of the CWG32 and CWG64-32 generators was $2^{37}$ and $2^{38}$ bits while they passed the tests up to $2^{43}$ and $2^{44}$ bits, they can generate numbers up to a threshold that is 64 times higher than is required by the minimum guaranteed period without incurring failure. The period of the CWG16 and CWG32-16 generators was $2^{20}$ and $2^{21}$ bits while they passed the tests up to $2^{28}$ and $2^{24}$ bits (which is consistent with the expected path to the cycle and the cycle length). Therefore, they can generate numbers up to a threshold 8-256 times higher than is required by the minimum guaranteed period without incurring failure. Hence, we expect that larger-scale versions of these generators would also pass the tests above the threshold of the minimum guaranteed period. NIST, Dieharder, and TestU01 were not applied to the 16-bit versions because they generated too small a number of bits to pass these tests.
@@ 7.3 Birthday Test
Additionally, we performed 1000 runs of the birthday test on a scaled-down version of the Collatz-Weyl Generator CWG32 and 32768 birthday tests on version CWG16 [27]. First we calculated how many outputs to look at and how many repeats we expect to see, according to an estimate of the expected number of repeats based on a binomial distribution:
$r\approx n-d\left(1-\left(1-\frac{1}{d}\right)^{n}\right)$
, where d is the number of possible different outputs that the generator can produce, and n is the number of samples; then we collected outputs, and counted how many repeats we saw. Finally we calculated p-values for the obtained results. A birthday test can cause some popular PRNGs to fail quickly because non-zero repetitions are expected in a true random number generator over its period of operation. On the contrary, the numbers generated by most PRNGs cannot be repeated in the range of their period. This is not the case with Collatz-Weyl Generators. The average p-value calculated based on observed repetitions in 1000 runs of the CWG32 version, initalized with random seed values, for the birthday test was $0.507$. In the case of CWG16 we tested all possible $2^{15}$ seeds $s$ (every odd $s$, with other variables set to 00) in the full range of minimum guaranteed period ($2^{16}$ outputs). In an ideal random number generator, the average expected number of duplicates in the range of $2^{16}$ 16-bit numbers is 24109.2, while the average number of duplicates in all streams of the CWG16 generator was 24108.4, and the average p-value was $0.500$. These results are therefore indistinguishable from true random number generation under the birthday test and confirm that the Collatz-Weyl Generators can be used in the full range of their minimum guaranteed period without incurring the risk of failing that test.
@@ 8 COMPARATIVE TESTS OF THROUGHPUT
We measured the time taken by the generators to generate $10^{9}$ random numbers by using an AMD Ryzen 5 4600H, 3 GHz processor with an Ubuntu Windows Subsystem for Linux (WSL) with gcc version 9.4.0. We also assessed the performance of several popular PRNGs for the sake of comparison. Ranlux++ and MixMax17 were run on a MacBook Pro 2.6 GHz, Intel Core i7 CPU in the ROOT library developed by CERN [28]. Table 3 lists the results.
@@ 9 CONCLUSION AND FUTURE RESEARCH
Let us consider the simplest Collatz-Weyl Generator with $m=2$ that is based on the original Collatz sequences, as shown in Figure 13 (or, equivalently, presented in Figure 14). This generator outputs only one bit per call, but it becomes very difficult for an attacker who does not know $s$, $weyl$, and other bits of $x$ to predict the next or past outputs of the generator after skipping the first few outputs. This is not due only to the non-invertibility of the mapping itself (to iterate back, we have to guess at least one bit at every step even if we know all the states of the generator). The least significant bits of such sequences generated by the proposed algorithm exhibit deterministic chaos, similar to that observed in Collatz sequences and their generalizations [29, 30, 31, 32, 33], and thus passed all stringent tests of randomness. This raises the prospect of the possibility of constructing efficient cryptographic primitives based on such schemes—for example, an ultra-lightweight stream cipher. Note that if we define the states of the proposed generator on registers of unlimited size, it seems we obtain a generator with an infinite period. However, we set aside this issue as a hypothesis.
@@
Acknowledgements
I would like to thank professor Dan Tamir, with whom I established collaboration in 2021 and whose article [31] inspired me to develop my ideas. During the two years of our collaboration, the schemes I worked on evolved and were significantly modified by me under the pressure of Dan’s critical comments.